please correct this problem If a winner had 220 more votes than the looser and the total votes were 810 how many votes did each have?
Answer by vasu2qute(17)

Let the loser's votes be = x
Then the winner's votes = x+220
Total votes = x + (x+220) =810
that is
x + x+ 220 = 810
2x+220 = 810
2x = 810 -220
2x = 590
x = 590/2
x= 295
x+220 = 295+ 220 = 515
Therefore the loser gets 295 votes
and the winner gets 515 votes